You are the Vice President for Human Resources at XYZ Investment Management, Inc. Your employees have recently complained about their long working hours. You collected a sample of 49 timesheets and calculated an average of 41.5 hours with a standard deviation of 2.5 hours. Using a significance level of 0.01, can you conclude that your employees work more than 40 hours per week?
Answer by stanbon(75887)

You collected a sample of 49 timesheets and calculated an average of 41.5 hours with a standard deviation of 2.5 hours. Using a significance level of 0.01, can you conclude that your employees work more than 40 hours per week?
---------------
Ho: mu = 40
Ha: mu > 40 (Claim)
----------------
Critical value for one-tail T-test with alpha=1% and df= 48 = 2.326
---------
Test statistic: t(41.5) = (41.5-40)/[2.5/sqrt(49)] = 4.2
-------------------
p-value = P(4.2 < t < 10)= 0.000057539...
---------------
p-value is significantly less than alph, so reject Ho.
The test supports acceptance of the Claim.
==============================
Cheers,
Stan H.